# NOBLE - Neural Operator with Biologically-informed Latent Embeddings to Capture Experimental Variability in Biological Neuron Models

### v1

@@ Abstract
2 Department of Computing and Mathematical Sciences, California Institute of Technology, Pasadena, CA, USA
3Department of Neurosurgery, Cedars-Sinai Medical Center, Los Angeles, CA, USA
4Department of Neurology, Cedars-Sinai Medical Center, Los Angeles, CA, USA
5Center for Biomedical Science, Cedars-Sinai Medical Center, Los Angeles, CA, USA
6Archimedes AI, Athena Research Center, Marousi, Greece
@@ Acknowledgements
L.G. was responsible for the complete technical implementation of this work. L.G., V.D., and B.T. jointly conceptualized, formalized, and developed the technical formulation of the novel NOBLE framework. P.H.W. supplied the biophysical PDE models, neuron morphologies, and neuroscience-specific domain expertise. L.G., V.D., B.T., and P.H.W. co-wrote the manuscript. C.A.A. and A.A. provided supervision and insightful editorial comments.
@@
A.A. is supported by the Bren Endowed Chair, ONR (MURI grant N00014-23-1-2654), and the AI2050 Senior Fellow program at Schmidt Sciences. C.A.A. and P.H.W. are being supported by the National Institutes of Health R01 - NS130126.

### v2

@@ Abstract
2Department of Computing and Mathematical Sciences, California Institute of Technology, Pasadena, CA, USA
3Department of Electrical and Computer Engineering, University of Alberta, Alberta, Canada
4Neuroscience and Mental Health Institute, University of Alberta, Alberta, Canada
5Department of Neurosurgery, Cedars-Sinai Medical Center, Los Angeles, CA, USA
6Department of Neurology, Cedars-Sinai Medical Center, Los Angeles, CA, USA
7Center for Biomedical Science, Cedars-Sinai Medical Center, Los Angeles, CA, USA
8Archimedes AI, Athena Research Center, Marousi, Greece
@@ Acknowledgements
L.G. was responsible for the complete technical implementation of this work. C.A.A. and A.A. conceptualized this work. L.G., V.D., and B.T. jointly developed the methodology of the novel NOBLE framework. P.H.W. and C.A.A. provided neuroscience-specific domain expertise, contextualizing the relevance and impact of this work within the broader field of neuroscience. P.H.W. supplied the biophysical PDE models and neuron morphologies. L.G. and P.H.W. produced the figures. L.G., V.D., B.T., P.H.W., and C.A.A. co-wrote the manuscript. C.A.A. and A.A. provided supervision and insightful editorial comments.
@@
A.A. is supported by the Bren Endowed Chair, ONR (MURI grant N00014-23-1-2654), and the AI2050 Senior Fellow program at Schmidt Sciences. C.A.A. is being supported by the National Institutes of Health R01 - NS120300 and R01 - NS130126. P.H.W. is being supported by the National Institutes of Health R01 - NS130126.